OpenStudy (anonymous):

I don't want the answer. I just want help working it out. What is the slope of y=3.5x

OpenStudy (anonymous):

you know: y=mx ok?

OpenStudy (ageta):

5, y from -10 to 10, 3.5x -1). 23.) Graph the linear equation x-2y=8 -2y=-x+8 y=-x/-2+8/-2 y=x/2-4

OpenStudy (anonymous):

according to the formula: we have: y=3.5x and m=3.5

OpenStudy (anonymous):

So, @felavin 3.5 is the slope?

OpenStudy (anonymous):

yep.

OpenStudy (anonymous):

ok, well @felavin can you help me with something else?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

I need to find the y-intercept as well. How do I find it? @felavin

OpenStudy (anonymous):

according to the: y-y0=m(x-x0)

OpenStudy (anonymous):

I don't understand that. @felavin

OpenStudy (anonymous):

if you get the y-intercept just need a m (slope) and one point.